Clindamycin 5 mg/kg/day IV QID for infection is ordered. you have Clindamycin 150 mg/ml. the child weighs 44lbs

a.how many mg per dose will the child receive?

b. How many mg per day will the child receive/

c How many ml per dose will you draw from the vial?

2. Amoxicillin 5 mg/kg/day divided q 6hrs is ordered for a child weighing 45lbs. Amoxicillin suspension is available in 10 mg/2ml

a how many mg per dose will the child receive

b. how many mg per day will the child receive

c how many ml per dose will the child receive

3. the order is for Rocephin 1.5 mg/kg IM one time dose. The infant weighs 24lbs. Rocephin comes in 5mg/5ml

a. how many mg per dose will the child receive

b how many ml per dose will the child receive

4. Convert the following

a 1L= ..........ML

B. 1kg=.......lb

c 15g+=........mg

d 1 cup=.....oz

e. 10mg=......mcg
Answer #1

Clindamycin order= 5 mg/kg/day IV QID for infection

Available clindamycin 150 mg/ml.

Child weight= 44lb=20kg

a.The mg per dose will the child receive= 5mg X 20= 100 mg per day and per dose = 100/4=25 mg as the order is Qid means four times a day.

b. How many mg per day will the child receive=100 mg

c How many ml per dose will you draw from the vial=?

                                                                                 = 25/150=0.16 ml

2. Ordered Amoxicillin= 5 mg/kg/day divided q 6hrs

Child weight= 45lbs= 20.45 kg

Available Amoxicillin suspension=10 mg/2ml

Means 5mg/ml

a how many mg per dose will the child receive=102.27/4=25.56 mg

b. how many mg per day will the child receive=5X20.45=102.27 mg

c how many ml per dose will the child receive= 25.56/5=5.11 ml

3. Ordered Rocephin= 1.5 mg/kg IM one time dose.

The infant weighs 24lbs=10.90 kg

Rocephin comes in 5mg/5ml= 1mg /ml

a. how many mg per dose will the child receive= 10.90X 1.5=16.36 mg

b how many ml per dose will the child receive= 16.36 ml

4. Convert the following

a 1L= 1000 ML

B. 1kg=2.2 lb

c 15g=15000mg

d 1 cup= 8 oz

e. 10mg=10000mcg